OpenStudy (anonymous):

Find the slope of the line that passes through (3, -4) and (-1, 2). A. 2/3 B. 3/2 C. -2/3 D. -3/2

OpenStudy (hoblos):

(2+4)/(-1-3) = 6/(-4) = -3/2